OpenStudy (anonymous):

Solve the proportion. 7/d + 28/d + 2

OpenStudy (anonymous):

CORRECTION: 7/d = 28/d +2

OpenStudy (anonymous):

7=28+2d d=-21/2

OpenStudy (anonymous):

\[7\div d = 28 \div d + 2\]

OpenStudy (anonymous):

You don't add d and 2 together. You cross multiply. But I don't understand where to go from there.

OpenStudy (anonymous):

\[\frac7d=\frac{28}d+2\] \[\frac7d=\frac{28+2d}{d}\] 7=28+2d

OpenStudy (anonymous):

Nevermind. got 2/3

OpenStudy (anonymous):

Thank you for helping though.

OpenStudy (anonymous):

wait wait...

OpenStudy (anonymous):

yes?

OpenStudy (anonymous):

your question is\[\frac7d=\frac{28}{d+2}\]

OpenStudy (anonymous):

Yeah!

OpenStudy (anonymous):

you posted it wrong

OpenStudy (anonymous):

it looks like 7/d=(28/d)+2

OpenStudy (anonymous):

however....now.

OpenStudy (anonymous):

I did. I didn't know how to put it in that form.

OpenStudy (anonymous):

7(d+2)=28d 7d+14=28d 21d=14 d=14/21=2/3

OpenStudy (anonymous):

i think now it is ok

OpenStudy (anonymous):

Yeah! Correct! haha.

OpenStudy (anonymous):

thanks